### v5

# ConstMig: Enabling Secure Live Migration of Large Intel SGX-based applications

Abstract.
Cloud service providers are adopting Trusted Execution Environments, or TEEs, to provide hardware-guaranteed security to applications running on remote, untrusted data centers. However, migrating such applications still uses the decade-old stop-and-copy-based method, which introduces large downtimes. Modern live-migration approaches such as pre-copy and post-copy do not work for such applications due to limitations imposed by the TEE.
We propose ConstMig, a near-zero downtime live-migration mechanism for large memory footprint TEE-based applications. ConstMig is fully compatible with containers, virtual machines (VMs) and microVMs. Our prototype, built on Intel SGX, a TEE solution from Intel, has a near-zero downtime irrespective of enclave size. ConstMig reduces the total downtime by 77-96% for a suite of SGX applications with multi-GB memory footprints compared to state-of-the-art TEE-based migration, MigSGX.
1. Introduction
Cloud service providers, or CSPs, are adopting trusted execution environments, or TEEs, to execute security-sensitive applications (Bacon et al., 2010; Orobosade et al., 2020; Singh and Chatterjee, 2017; Baumann et al., 2014; Gribov et al., 2017; Wang et al., 2020; Priebe et al., 2018; Antonopoulos et al., 2020; Dahl et al., 2018; Hong et al., 2020; tf-encrypted, 2025; Tramer and Boneh, 2019; Shafran et al., 2021) on remote, untrusted cloud nodes. In a TEE setting, the hardware ensures the application’s security. The hypervisor and operating system (OS) are assumed to be untrusted (§2.1). Leading CSPs such as Microsoft and Google provide (Cfir Cohen, 2023; ju-shim, 2023) several TEE options such as Intel SGX (Intel, 2019c), Intel TDX (Cheng et al., 2024a), AMD SEV (Kaplan et al., 2016), ARM TrustZone (Alves and Felton, 2004) (with Confidential Compute Architecture or CCA planned (Ltd., )), and RISC-V Keystone (Lee et al., 2020).
Figure 1. ConstMig ensures a non-zero total downtime while migrating an SGX-based application, irrespective of its size. TDX enables a hardware-based live migration but has long boot/respond time.
Deploying TEEs in data centers presents unique challenges in migration of applications from one machine to another. CSPs rely on efficient live migration to provide high availability, load balancing, efficient resource management, failure tolerance, seamless upgradation, and code patching (Khanna et al., 2006; Arzuaga and Kaeli, 2010; Zaw and Thein, 2012; Jin et al., 2011). In a non-TEE setting, CSPs rely on modern live migration mechanisms such as pre-copy (Clark et al., 2005) and post-copy (Hines et al., 2009) employed by hypervisor or a migration manager (§2.4) to achieve a near-zero downtime during migration (Clark et al., 2005; Hines et al., 2009). However, these techniques are not available in a TEE setting as they require access to either userspace fault handling (userfaultfd) or tracking page writes (using DIRTY bits in the page table) which are prohibited by the TEE hardware (see Section 2.4(Li and Li, 2024; Labs, ; Wang et al., 2022) as the operating system (OS), hypervisor, or any other migration manager is not a part of the trusted computing base (TCB).
Recognizing this challenge, TEE vendors integrated hardware support to enable live migrations of TEE applications. A migration manager typically orchestrates the migration and uses the API exposed by the hardware to handle the secure data pages to TEE applications – which otherwise would be inaccessible (or ineligible) to the migration manager due to TEE security guarantees. AMD enables migration of secure VMs using Migration Agents – special VMs that provide migration keys, enforce guest policies, and form part of the VM’s TCB (61). Intel TDX uses a small VM known as MigTD that assists the untrusted VMM in accessing the secure state (after encryption). ARM TrustZone does not support live migration; however, the proposed ARM CCA will have hardware-enabled live migration (Ltd., ). Intel SGX, which allows for TEE enablement at the per-application level, has no hardware support for migration. Hence, prior work in enabling migration for SGX relied on the classical stop-and-copy method for migrating SGX enclaves (Nakashima and Kourai, 2021; Liang et al., 2019; Alder et al., 2018; Guerreiro et al., 2020; Soriente et al., 2019). In the stop-and-copy method, an executing application is paused, a checkpoint is taken, and the checkpoint is transferred to the destination machine, where it is subsequently resumed. This approach has a large downtime, which is proportional to the memory footprint (see Figure 1). Given that Intel has planned support for SGX till Intel Xeon 7 – Diamond Rapids(Larabel, ), we believe that missing live migration support is a key omission which this paper aims to solve. Secure VMs provided by Intel TDX, which do have hardware support for live migration, suffer from malicious co-hosted applications, a large TCB, and a high memory encryption cost even in the unsecure portions of an application (performance overheads). SGX has a small TCB, low initialization time, and no overheads for the unsecure portion (§2.3).
We present ConstMig, a novel live migration mechanism for SGX enclaves with a near-zero downtime that is independent of the memory footprint of the application. ConstMig requires no additional hardware and has far lower performance overheads than the state-of-the-art (see Figure 1). ConstMig addresses a fundamental challenge in enabling live migration for enclave-based systems: a fine-grained tracking of memory accesses on the destination machine to ensure that the restored application reads valid data from the source machine – instead of an invalid, all-zero page (§4). We propose two access tracking mechanisms: ❶ based on fine-grained page fault tracking ( §4.1) using kernel modifications and ❷ based on fine-grained pointer access tracking (§4.2). In addition to providing all the security guarantees of SGX, ConstMig also provides a freshness guarantee during the migration process, which ensures that an attacker cannot replay old encrypted enclave data during migration (§5).
Our key contributions are as follows:
  1. (1)
    We identify the key challenges faced while adopting modern migration schemes for enclaves.
  2. (2)
    We propose a novel migration scheme that addresses these challenges and enables near-zero downtime while migrating large-sized SGX applications. To our knowledge, we are the first to implement this and show its benefits.
  3. (3)
    We propose novel optimizations during the migration process that reduce the total migration time and reduce the cost of migration.
  4. (4)
    We evaluate our solution using a set of widely used SGX benchmarks. Our scheme outperforms the current state-of-the-art (Nakashima and Kourai, 2021) by 77-96% for a suite of Intel SGX applications that have multi-GB memory footprints.
2. Background & Motivation
Here, we discuss some relevant background for the paper.
2.1. Intel Software Guard eXtensions or SGX
Intel Software Guard Extensions (SGX) enable secure execution of applications on untrusted machines by creating a secure sandbox called an enclave (Costan and Devadas, 2016). These enclaves are isolated from even privileged entities like the OS, hypervisor, and system administrators. At boot, SGX reserves a portion of main memory which is hardware-managed and encrypted using a hardware key inaccessible to any software component. The usable part of this reserved space is called the Enclave Page Cache or EPC (the rest is used by SGX to store metadata). The latest version of SGX, referred to as scalable SGX, supports an EPC of up to 512 GB (Intel, 2022b, a), compared to the earlier version (SGXv1) with a usable EPC of only 92 MB (Kumar et al., 2022a; Kumar and Sarangi, 2021; Intel, 2019a).
Memory management: SGX ensures an enclave’s integrity by validating the signature of a loaded enclave (Costan and Devadas, 2016) at initialization time. However, this process can become costly for enclaves with large memory footprints, as all pages must be provisioned upfront. For instance, as per our experiments, initializing a 20 GB enclave takes $\approx 30$ seconds. To address this, Intel introduced Enclave Dynamic Memory Management (EDMM) (McKeen et al., 2016), which allows developers to specify a minimal set of committed pages during initialization, while additional uncommitted pages can be added dynamically (Kumar et al., 2025). In principle, EDMM enables an enclave requiring 512 GB to start with a single committed 4 KB page, though this flexibility may introduce performance overhead depending on the workload (Kumar et al., 2025; Dhanraj et al., 2025).
extitSGX execution framework: SGX provides security at the application level. Here, an application is partitioned into secure and unsecure components. The secure component executes within an enclave, incurring performance overheads due to security guarantees, whereas the unsecure component runs conventionally without overhead (see Figure LABEL:fig:scoping). Interaction between the two occurs via enclave calls (ECALL) from the unsecure side and outside calls (OCALL) from the enclave (Costan and Devadas, 2016).
2.2. Intel Trusted Domain Extension or TDX
Intel Trusted Domain eXtension or TDX is the next-generation TEE offering from Intel which introduces the concept of a hardware-protected confidential virtual machine (CVM) (Misono et al., 2024; Kuvaiskii et al., 2024)– also known as Trusted Domain (or TD)– instead of just a single application. TDX allows unmodified applications to execute inside a hardware-protected VM (although the OS and the hypervisor must be modified to handle TDX specific interrupts (Cheng et al., 2024b; Misono et al., 2024)). TDX enables live migration of a CVM with hardware support.
Intel TDX supports dynamic memory management through the Dynamic Physical Address Metadata Table (PAMT), which stores metadata essential for securely managing physical memory assigned to Trust Domains (TDs). While the entire DRAM is encrypted using Intel Multi-Key Total Memory Encryption (MKTME), TDX enables dynamic allocation and deallocation of PAMT entries, allowing memory pages to be added to or removed from a TD at runtime, enabling full DRAM encryption and integrity protection without huge initialization overheads.
2.3. Intel SGX vs Intel TDX
Intel TDX is the latest TEE offering from Intel which aims to secure the whole VM and has hardware support for a live migration. TDX is the natural choice for deploying an application with TEE support with minimal development effort. However, SGX brings its own advantages that are either lacking or missing in TDX.
2.3.1. Execution Overhead
As discussed, Intel SGX divides an application into secure and unsecure portions. The secure portion executes in an enclave inside SGX and incurs the overhead due to security guarantees of SGX. The untrusted portion, however, suffers no additional performance overhead. However, Intel TDX does not follow such a partitioned approach. An application executes completely within a CVM and all of its data is protected by the TDX module (Misono et al., 2024).
[…]
CVM switch overhead: In a TEE solution, frequent transitions between secure and unsecure worlds and memory encryption are the main causes of performance overheads. Misono et al.(Misono et al., 2024) show that workloads that require frequent VM exits can suffer performance slowdowns of up to 240% and 472% on AMD SEV and Intel TDX, respectively, compared to traditional VMs. The additional overhead in TDX is mainly due to the involvement of the TDX module in handling VM exits, which adds extra processing steps (Misono et al., 2024).
extbfSGX Advantage: SGX’s flexible execution framework can address the performance concerns due to security measures. With SGX, the performance overhead in terms of memory encryption and costly enclave entry and exit are relevant only for the secure portion of the application (§2.1). The untrusted part executes in a conventional manner with no performance loss. Figure 2 shows the end-to-end performance of Key-Value workload for varying amount of sensitive data – for 100% everything executes in SGX and, say, for 0% only 60% data is in SGX, rest in the untrusted region outside SGX. With 100% of data as secure, SGX has a performance overhead of $\approx$12% compared to TDX. However, with 20% of data as secure, SGX outperforms TDX by $\approx$ 9% and approaches the performance of No-SGX.
Figure 2. Performance of Key-Value with No-SGX, TDX, and SGX for varying amount of secure portion of the allocated data.
2.3.2. TCB Size:
The TCB (Trusted Computing Base) of TDX includes the TDX-enabled processor with Virtualization Technology (VT), Multi-key Total Memory Encryption (MKTME) and SGX (Cheng et al., 2024b), along with TDX Module ($\approx$20 K LOC (Intel, 2024)), and the entire guest OS and software stacks within the CVM (Cheng et al., 2024b). Such a large TCB provides a wide attack surface and makes Intel TDX inherently susceptible to errors and vulnerabilities(Misono et al., 2024). Misono et al.(Misono et al., 2024) report that 54% of the identified vulnerabilities related to Intel TDX are associated with the underlying firmware. SGX’s TCB consists of the processor, the microcode and firmware that provide security, and the secure part of an application’s code. The untrusted portion of the application and the host operating system are not part of the TCB.
2.3.3. Initialization overhead
After the network transfer time, the time to initialize an enclave/CVM affects the migration time the most ($17\%$, see Figure 3). With SGX, the initialization time of an enclave is proportional to the number of committed pages – a low count of committed pages reduces the initialization time, but may incur performance overhead (Kumar et al., 2025; Dhanraj et al., 2025). TDX uses PAMT to dynamically add 4 KB pages after booting. We observed that a TDX CVM on an Emerald Rapid Machine (§6.1) takes $\approx 26$ seconds to boot up, irrespective of the allocated memory (similar observation made by Misono et al.(Misono et al., 2024)).
2.4. Migration methods and Challenges with SGX
The performance of a migration method is evaluated using two metrics: total downtime – the period during which the application is inactive – and total migration time – the duration from migration start to completion (Stoyanov and Kollingbaum, 2018; Nakashima and Kourai, 2021).
Figure 3. For stop-and-copy, for large enclaves, the total migration time is dominated by the enclave initialization time on the destination machine and the total transfer time.
2.4.1. Stop-and-Copy Migration
In the stop-and-copy migration approach (Clark et al., 2005), the application is paused on the source machine while its state (e.g., registers, memory, open files, network) is transferred to the destination machine. It resumes execution only after the transfer completes. As a result, both total migration time and downtime scale are proportional to the application’s memory footprint.
[…]
Figure 3 shows the four phases of stop-and-copy migration: save, transfer, initialize and restore. For small enclave sizes ($\approx$ 10 MB), the total migration time is dominated by the save and restore phases. However, for larger enclave sizes (>>100 MB), the network transfer phase dominates the total migration time (up to 96% for 10 GB+ enclaves).
Figure 4. pre-copy and post-copy migration mechanisms. The total downtime is independent of the memory footprint. Start,Pause, and Resume indicate the start of migration, pausing, and resuming of the application, respectively.
2.4.2. Pre-copy Migration
As shown in Figure LABEL:fig:pre_copy, pre-copy migration proceeds as follows: ❶ Snapshot the application’s state and track subsequent write operations. ❷ Transfer the snapshot to the destination. ❸ Identify and track modified (dirty) pages. If their count is below a threshold, ❹ pause the application and transfer the dirty pages else send the dirty pages again and go back to step three. ❺ Resume the application on the destination. Here, total migration time scales with memory footprint, while downtime depends on the number of dirty pages in the final transfer.
Challenge in adopting for SGX: Tracking an application’s address space to identify dirty pages is essential for the pre-copy migration method. Non-TEE migration approaches typically rely on the page table entry’s dirty bit, which is automatically set by the hardware upon a write to a data page (64; 66; 27). Prior work (Bulck et al., 2017b, a) has shown that dirty bits are also set for SGX-backed pages, enabling the tracking of write operations within enclaves. However, these findings were based on experiments conducted on the Intel Skylake i7-6700, a desktop-class processor released almost a decade ago (Intel, 2025a). Since then, Intel has discontinued SGX support on desktop CPUs, restricting it to server-class processors (to Wikimedia projects, 2015), and has introduced several security enhancements to mitigate SGX-related attacks (Randmets, 2021; Fei et al., 2021; Intel, 2025c).
In our experiments using the latest Xeon-based servers (Icelake and Emerald-Rapids), we observed that dirty bits are not set for SGX-backed pages, even after repeated write operations (see Appendix A for a micro-benchmark to validate this claim).
2.4.3. Post-Copy Migration
In post-copy migration (Figure LABEL:fig:post_copy), ❶ the application is first paused on the source machine, and a minimal state (CPU registers, non-pageable memory (Hines et al., 2009; Clark et al., 2005)) is sent to the destination machine. ❷ The application is then resumed on the destination, while the remaining state is transferred in the background. ❸ During the transfer, memory accesses to pages still on the source machine trigger network page faults, which are transparently handled by the OS and migration mechanism. Here, total migration time scales with memory footprint, but downtime remains constant. However, performance may degrade during migration due to costly network faults.
Challenge in adopting for SGX: Tracking memory accesses of a restored application is critical for the post-copy migration. In non-TEE environments, migration frameworks use the userfaultfd mechanism (Doc, 2022) to intercept page faults. With userfaultfd, the OS forwards page faults to the migration manager, which then fetches the missing page from the source machine and services the fault. However, userfaultfd supports only anonymous memory, huge pages, and shared memory regions (Doc, 2022). As SGX enclave memory is device-backed via /dev/isgx (Corbet, 2023, 2022; Doc, 2022; Kernel, 2023), userfaultfd cannot track memory accesses within enclaves.
Figure 5. ConstMig handles the migration of the secure portion of an application. The rest is handled by CRIU.
3. Design
In this section, we discuss the design of ConstMig. As shown in Figure LABEL:fig:scoping, ConstMig handles the migration of the secure portion of the application. CRIU handles the migration of the untrusted portion. The primary design goals of ConstMig are as follows:
  1. (1)
    Low downtime: Enable live migration for large SGX enclaves with minimal total downtime
  2. (2)
    Performance overhead: Minimize performance degradation due to migration.
  3. (3)
    Security: Maintain confidentiality, freshness, and integrity of the data at all times during migration.
We maintain the following invariants during the migration operation:
Invariant 1: It is never possible that a page fault with virtual address $V$ on the destination machine is served with the data of a page with virtual address $V^{\prime}$ ($V\neq V^{\prime}$).
Invariant 2: Once the migration key is transferred, the source cannot be resumed.
Invariant 3: A page is “saved” and “restored”, at most once.
Invariant 4: Migration components executing outside SGX never see the plain text data.
Invariant 5: The migration key ($M_{k}$) is a function of the PUF of the source machine, boot time, current time, and the output of a TRNG. A replay attack is thus not possible (§ 5).
3.1. Scope
As discussed earlier, modern TEEs support hardware-assisted live migration, with the notable exception of Intel SGX. Given Intel’s plans to continue SGX support on future server-class processors, enabling live migration for large SGX enclaves remains an important and timely challenge. Our focus is primarily on heap memory, which constitutes the majority of an application’s memory footprint (Aga and Narayanasamy, 2019). In contrast, other segments such as data and BSS typically range from a few KBs to MBs and are always transferred synchronously as part of the minimal state during migration.
Furthermore, we do not discuss migration of sealed data. SGX provides a way to seal data using a hardware-bound key. Alder et al.(Alder et al., 2018) proposed a method to migrate sealed data by encrypting all the sealed data with a software-generated key on the source machine and resealing it on the destination machine after transfer and decryption. This solution can be easily incorporated into ConstMig.
3.2. Threat Model
We adopt the same security assumptions as Intel SGX as prior work in this domain. The operating system and other software components outside the enclave are not trusted. They do not have access to enclave plaintext data or the cryptographic keys used during migration.
In SGX, the OS manages the page tables of secure applications. A malicious OS can manipulate the PRESENT bit in page table entries (PTEs) to infer the application’s page access patterns (Xu et al., 2015). Such side-channel attacks fall outside the threat model of SGX and are beyond the scope of this work (see §5). Migration is aborted immediately upon detection of any security violation.
3.3. High-Level Overview
There are two main components in ConstMig: the migration manager (CRIU) that orchestrates the entire migration process and a trusted SGX library (ConstMigLib), which is linked to the secure application (see Figure LABEL:fig:scoping). Its routines have access to the secure memory space of the application running in the enclave.
Start migration: As shown in Figure LABEL:fig:optmig_high_level, at the beginning of the migration, ❶ the migration manager pauses the application and sends the save signal to ConstMigLib. The migration manager then resumes a single thread for ConstMigLib to do its work (savethreadsave\,\,thread). ❷ ConstMigLib generates a master key $M_{k}$ to be used during migration. It then saves the data segment, BSS segment and OptMgr metadata (information about the allocated regions) in an untrusted metadata buffer (MBuff). ❸ A thread starts to save heap data in a separate untrusted backup buffer (BBuff). All these operations are done after performing encryptions and adding integrity checks (§3.5.2 and §3.5.3).
Handshake: ❹ The migration manager sends a signal to spawn a fresh enclave on the destination machine after the start of the migration operation. Once the enclave is initialized, ConstMig first performs a remote attestation (Anati et al., 2013) to verify that only a valid enclave with ConstMigLib is running on the machine, thereby establishing the “chain of trust” on the destination machine. Once the enclave is validated, ConstMig transfers $M_{k}$ to the destination machine using a secure key transfer protocol that ensures that the key is only transferred to the intended destination machine (Anati et al., 2013). The original enclave cannot be resumed after this (see Invariant 2). If migration fails for any reason, the application instance is lost on both machines. ❺ Following this, the destination machine receives the encrypted MBuff.
Start Restore: ❻ ConstMigLib on the destination machine, decrypts MBuff using $M_{k}$ and reads the encrypted BSS segment, data segment, and memory manager’s metadata. ❻ The migration manager on the destination machine sends the restore signal to ConstMigLib. ❼ A background thread starts reading data from BBuff and starts restoring the enclave’s heap (restorethreadrestore\,\,thread). Since BBuff is stored in the untrusted region, it is protected by userfaultfd (Doc, 2022), and the restore thread is guaranteed to read valid pages.
Cleanup: Once the restoration is done, ConstMigLib zeros all the allocated pages on the source machine and kills the application – preventing forking attacks (§5). At any given time, an application runs on only one system.
3.4. Implementation Details
3.4.1. Migration Manager
We use the widely popular Checkpoint/Restore In Userspace or CRIU (CRIU Project, 2025; Venkatesh et al., 2019; Stoyanov and Kollingbaum, 2018; Nakashima and Kourai, 2021) package as our migration manager. CRIU supports both the pre-copy and post-copy live migration methods. In ConstMig, CRIU is not a part of the trusted code base, and never sees the migrating enclave data in plaintext.
To support SGX enclave migration, we extend CRIU with enclave-aware functionality. A save signal initiates a state save process and marks the beginning of a migration, while a restore signal triggers restoration when resumed on the destination machine. The signals are captured and processed by ConstMigLib.
extbfEncrypted buffers: We utilize two encrypted buffers during migration: a metadata buffer (MBuff) that stores the BSS, data segments, and ConstMig-specific metadata, and a backup buffer (BBuff) to store the encrypted heap contents. Both buffers are allocated in untrusted memory outside the SGX enclave.
Due to its relatively small size, MBuff is transferred synchronously to the destination at the beginning of migration. In contrast, BBuff – containing large encrypted heap data – is transferred asynchronously. We integrate the userfaultfd mechanism via CRIU to support on-demand page retrieval for BBuff. Since it resides in untrusted memory, read operations on the destination machine either access already-transferred data or trigger synchronous page fetches over the network (Hines et al., 2009).
A consistency checker: CRIU handles the restoration of BBuff on the destination machine. However, CRIU is not SGX migration aware. To ensure correctness, we enforce two key constraints:
  1. (1)
    BBuff is transferred only after the enclave state has been successfully saved on it.
  2. (2)
    The restored enclave only reads restored data.
To enforce these constraints, our consistency checker maintains three bit-vectors—internal save_vec, save_vec, and restore_vec of size equal to the number of 4 KB pages in BBuff. The internal save_vec, allocated within the trusted region, tracks when a page is ready for transfer. A copy of this vector, save_vec, is maintained in the untrusted region for CRIU’s use. restore_vec, stored inside the enclave, indicates whether a page has been restored into enclave memory. All entries in save_vec and restore_vec are initialized to zero.
extitFine-grained Access tracking: To ensure that a restored enclave accesses only valid pages, we enforce that a page read is permitted only if the corresponding bit in restore_vec is set. Otherwise, the access is paused until the page is fetched from the source machine. Such support is missing from SGX as of now. We introduce two novel mechanisms to capture enclave memory accesses: access tracker and fault tracker. The access tracker automatically inserts memory access checks into the application’s source code (using an LLVM pass), enabling fine-grained control. In contrast, the fault tracker leverages page table entries to detect access attempts at runtime.
Note that using page table entries to track leaks the memory access pattern. However, as discussed earlier (§3.2), such leakage is unavoidable in SGX, as the OS –an untrusted entity –manages the page tables (§3.2). We provide a detailed comparison of these approaches in Section 4.
3.5. ConstMig library (ConstMigLib)
(a) Asynchronous data saving.
(b) Asynchronous data restore.
Figure 6. Asynchronous data restoring. Note that restore_vec is allocated inside an enclave as there is no need on the destination machine for CRIU to read it.
To perform migration, ConstMig requires knowledge of the application’s allocated memory regions, along with mechanisms to save and restore them. We achieve this by statically linking ConstMigLib to the SGX application, enabling support for memory region discovery and migration. ConstMigLib executes inside SGX and has access to all of the application’s address space in plain text form. Its main components are:
3.5.1. Memory Manager
We implement a custom memory manager (OptMgr) by overriding the default malloc() and free() functions to intercept all allocation and deallocation requests. For each allocation, we record the address, size, and a unique ID in a structured array (MemArr). Upon deallocation, the corresponding entry is removed from MemArr. During migration, MemArr is encrypted and stored in MBuff, which is transferred synchronously along with the enclave’s data and BSS segments to the destination machine.
3.5.2. Save thread
A dedicated save thread (save_th) handles the save command initiated by CRIU and begins capturing the heap contents into BBuff. As shown in Figure 6(a), operating at a granularity of 4 KB pages, save_th performs the following steps for each page $P$ at virtual address $VA_{p}$: it concatenates the page contents $PC(P)$ with the address to form $PC(P)||VA_{p}$, and computes a hash $H$ over this concatenation. The page contents and hash are then encrypted using $M_{k}$, and the resulting encrypted packet is stored in BBuff. Once a page is saved, the corresponding bits in both save_vec and internal save_vec are updated to indicate readiness for transfer to the destination machine.
3.5.3. Restore thread
A dedicated restore thread (restore_th) handles the restore command initiated by CRIU and begins restoring heap contents from BBuff. Operating at a granularity of 4 KB pages, the thread retrieves encrypted data from BBuff, decrypts it within SGX using $M_{k}$, and extracts the page contents ($PC(P)$) and hash ($H$). It then computes the expected hash ($H^{\prime}$) and verifies that $H==H^{\prime}$. If validation succeeds, the page is loaded into enclave memory; otherwise, a security violation is reported and migration is aborted (see Figure 6(b)). After successful restoration, restore_th sets the corresponding bit in restore_vec, marking the page as restored. If the bit was already set, indicating a duplicate restoration attempt, a security violation is reported and migration is aborted (see Invariant 3).
Network Faults: During restore, if the requested encrypted page is not available in BBuff (yet to be copied from the source machine), CRIU issues a network page fault to the source machine, and restore_th enters a blocked state. Upon receiving the request, the source machine’s CRIU signals save_th to save the page into BBuff. save_th first checks the internal save_vec to verify whether the page has already been saved. If so, a security violation is reported. Otherwise, the page is encrypted and stored in BBuff3.5.2), and CRIU transfers the encrypted page to the destination. Upon receipt, restore_th validates the encrypted contents before restoring them into enclave memory.
Algorithm 1 Prepare Encrypt Packet
1:procedure PrepareEncryptPacket($P,VA_{P},M_{k}$)
2:  Check(internal save_vec)
3:  $H\leftarrow\text{Hash}(PC(P)\|VA_{P})$ $\triangleright$ Link page content with address
4:  $C\leftarrow\text{Encrypt}(PC(P)\|H,M_{k})$
5:  Update (save_vec, internal save_vec)
6:  return $C$
7:end procedure
Algorithm 2 Validate Encrypt Packet
1:procedure ValidateEncryptPacket($C,VA,M_{k}$)
2:  Check(restore_vec)
3:  $PC(P)\|H\leftarrow\text{Decrypt}(C,M_{k})$
4:  $H^{\prime}\leftarrow\text{Hash}(PC(P)\|VA)$
5:  if $H=H^{\prime}$ then
6:   Update(restore_vec)
7:   return true
8:  else
9:   return false
10:  end if
11:end procedure
4. Fine-Grained Access Tracking
A key challenge in enclave migration is that both the restored enclave and SGX hardware are not migration-aware. Upon restoration, the enclave can access any addressable memory region; however, if it touches a page not yet restored by restore_th, it triggers a page fault. Instead of retrieving the original encrypted page from BBuff, the fault handler allocates a fresh zeroed page – leading to a use-before-restore issue. As discussed in Section 2.4, this cannot be mitigated using userfaultfd, since SGX pages are device-backed (via /dev/isgx/dev/isgx) and unsupported by userfaultfd in the Linux kernel (Corbet, 2022, 2023; Kernel, ; Kernel, 2023).
To address this issue, we propose two novel schemes: the fault tracker and the ConstMig access tracker. They ensure that a restored enclave reads valid data instead of a fresh all-zero page.
4.1. Fault Tracker
Figure 7. Working of Fault Tracker to ensure a consistent memory access in the restored enclave
The first approach (see Figure 7) modifies the SGX page fault handler to return restored data instead of zeroed pages on the destination machine. Since enclave page tables are managed by the OS, a fault is triggered when the enclave accesses an unallocated page. By default, the SGX fault handler returns a zeroed page; we extend it to fetch the corresponding encrypted page from BBuff (which may cause a network fault that is transparently handled by CRIU). However, the page must be decrypted before the application can use it. As the fault handler operates in kernel space and cannot decrypt enclave data, it invokes enclave services to perform decryption, validation, and placement of the restored page.
Fault capture and restoration: As illustrated in Figure 7, we extend the SGX driver to signal the application upon a page fault. Upon receiving the signal, the application performs an ECALL into the ConstMigLib enclave with the faulting address. ConstMigLib first checks whether the fault is due to an ongoing restoration by restore_th. If not, it retrieves the encrypted page from BBuff, decrypts and verifies its integrity, and writes the restored data to the faulting address. This ensures that the faulting thread receives the correct data instead of a zeroed page. Importantly, each enclave page triggers exactly one signal – either during restoration or due to a fault – maintaining Invariant 3.
Implementation: To intercept page accesses in the restored enclave, we leverage protection bits in page table entries (Gorman, 2012), setting PROT_NONE on committed pages via mprotect (68). We modify the sig_handler in the trusted SGX SDK (PSW (Intel, 2019a)) to handle protection faults and signal the application after clearing the protection bit. A dedicated trusted thread, constmig_fault_thread, is introduced in the SGX SDK to handle these signals. This thread performs an ECALL into ConstMigLib to decrypt, verify, and restore the faulting page. A separate thread is necessary since the faulting thread remains blocked until the fault is resolved; reusing it would crash the enclave.
4.2. Access Tracker
Figure 8. Working of Access Tracker to ensure a consistent memory access in the restored enclave
The second approach introduces a runtime check before each heap pointer access in the restored enclave (see Figure 8). During an active migration, this check consults restore_vec to verify whether the target pages have been restored from BBuff. Listing 1 shows an example using the __process_access() function, which executes entirely within the enclave and does not require an ECALL. Once migration is complete (all the pages are restored – indicated by a global flag), the memory check is replaced with return 0;.
The function __process_access() takes a pointer address and access size, performs a linear scan of MemArr– the array that contains information about all the allocated regions – to locate the corresponding memory region, and identifies the required pages in BBuff. While linear scans suffice for a small number of regions, a radix tree can optimize lookups at scale. It then checks restore_vec to determine if all required pages are already restored. If not, it reads the encrypted pages from BBuff (may cause network faults, handled transparently by CRIU), decrypts, and restores them.
1void traverseList(){
2 struct Node* pos = head;
3 __process_access(pos, sizeof(struct Node));
4 while(pos!=NULL){
5 __process_access(pos, sizeof(struct Node));
6 // use pos->val
7 pos = pos->next;
8 }
9}
Listing 1: List traversal showing the added heap pointer access checks
extbfImplementation: We implement a Clang-based source-to-source transformer (12; 67) to insert memory access checks before heap pointer dereferences. This technique, widely used for pointer safety (Nagarakatte et al., 2009; Richardson, 2020; Nagarakatte et al., 2015; Saileshwar et al., 2022), ensures that accessed pages have been restored. Our transformer extends Clang’s RecursiveASTVisitor to traverse the abstract syntax tree (AST) (1), which represents source code constructs such as variables and functions. For instance, the heap variable pos in Listing 1 appears as a DeclRefExpr node, containing its name and type. Using this information, the transformer inserts a call to __process_access() before each heap pointer access.
4.3. Fault Tracker v.s. Access Tracker
The choice between the two trackers depends on the extent of changes a cloud vendor is able to implement in its system.
Access to Source code Fault Tracker requires no enclave source code modification (apart from adding a signal handler in the untrusted part of the application). On the other hand, Access Tracker requires enclave source code modification to add checks before pointer accesses.
Page restoration Fault tracker restoration is expensive because it involves sending a signal and making an ECALL for each page restoration, whereas Access tracker is less overhead-prone because it does not require any signals or ECALLs (as the checks are done inside the enclave).
Figure 9 shows the time taken by each step in Fault and Access Trackers (resp.) for 4 KB data and 4 MB data, respectively. For 4 KB and 4 MB of data, Access Tracker outperforms Fault Tracker by $35\%$ and $93\%$, respectively. The performance improvement for Access Tracker grows with an increase in the size of the data to be restored. Fault Tracker operates at a 4 KB granularity; hence, restoring 4 MB of data will require $\approx$ 1,024 faults, signals, and ECALLs. Whereas Access Tracker can potentially use the MemArr to restore the complete 4 MB region using a single ECALL. The cost to restore a page (read from BBuff and decryption) is the same for both approaches. Since Access Tracker outperforms Fault Tracker, we use it for the rest of the experiments.
Figure 9. Performance of access tracker and fault tracker approach for consistency checker for 4 KB and 4 MB data. As fault tracker needs to fault on each page to restore it, the overhead of ensuring consistency checker is much higher compared to access tracker. FT: Fault Tracker, AT: Access Tracker
Post Migration After migration, Access Tracker has a small runtime overhead, which Fault Tracker does not have (§6.4).
5. Security Analysis
Table 1 lists the key components of ConstMig along with their trust level and the data or access pattern accessible to them. The design and implementation of ConstMig ensures protection against the following attacks.
Replay attack: An adversary might intercept messages transmitted during migration and attempt to replay them to migrate the application to a compromised destination. Any attempt to replay stale data will result in an integrity check failure during restoration (§3.5.3) and will lead to an immediate abort of the migration process (see Invariant 4). Replaying the data during the same migration session will result in the restore thread dropping the packet after decryption, as the restoration of the replayed page is already done (see Invariant 3).
Entity Trusted? Access
Operating system Page fault sequence
Migration manager (CRIU) Page fault sequence
Trusted Library (ConstMigLib) Everything
Save thread (save_th) Everything
Restore thread (restore_th) Everything
Table 1. Entities involved in migration and their trust level.
Forking attack: In a forking attack, the application is maliciously forked during a migration. Since $M_{k}$ is delivered only once, the forked instance will be unable to decrypt the pages, effectively preventing the attack. Moreover, ConstMig does not allow an application to resume on the source machine after the key transfer, preventing re-migration to any other destination (see Invariant 2).
Changing the Fault Address: An attempt by a malicious OS to disrupt the execution flow by changing the faulting virtual address from $VA_{A}$ to $VA_{B}$ will not result in any security leak. Based on the manipulated faulting request, the contents of the page $VA_{B}$ will be retrieved from the source machine in an encrypted form, which is guaranteed to be genuine (Invariant 1). Since the encrypted packet contains the address $VA_{B}$6(a)), the destination will write the page contents to address $V_{B}$, which is correct. Of course, the application will again immediately fault as the page fault for the address $VA_{A}$ is yet to be serviced. However, the point to note is that there is no security violation.
5.1. Fault Tolerance
Fault Tolerance A migration process may fail for several reasons: source crashes, destination crashes, network failures, etc. What happens in such a situation depends on the security guarantees of the environment. To handle such cases, we extend the proposal by  (Gu et al., 2017). The original enclave (OE) and restored enclave (RE) can communicate securely via remote attestation. ❶ RE, on receiving $M_{k}$, sends a message to OE. The save_th requires the message to start saving encrypted data in BBuff. If OE does not receive the message, it times out and resumes. Even if RE is executing, it will be a fresh instance of the enclave as it will not receive any data from OE. ❷ If it does receive the message, it is aware of two things: RE has received $M_{k}$ and has resumed. OE cannot be resumed now. If, for any reason, RE crashes, the application instance is lost. ❸ Once save_th is done, it zeros all the secure pages and kills OE (since all the data has been saved).
6. Evaluation
We evaluate two configurations of ConstMig: ConstMig-V1, where the enclave is initialized with the maximum required memory upfront, and ConstMig-V2, which leverages SGX’s EDMM capabilities to initialize with a smaller committed memory footprint and dynamically add pages as needed (Dhanraj et al., 2025; Kumar et al., 2025). We compare ConstMig with the current state-of-the-art MigSGX (Nakashima and Kourai, 2021). All other prior work (Nakatsuka et al., 2022; Liang et al., 2019; Guerreiro et al., 2020; Soriente et al., 2019) is based on a similar migration scheme. We also evaluate TDX– TDX with no migration due to the non-availability of two machines with TDX capability. We only report the initialization time for a CVM on the destination machine – a lower bound on the total downtime, as there will be other costs, such as save, transfer, and page faults (Intel, 2025b).
extitWorkloads: Table 2 lists the workloads used for evaluation. Most of the workloads in our evaluation setup either do not issue a lot of free() calls and hence are unaffected by the ratio of committed to uncommitted memory (Kumar et al., 2025; Dhanraj et al., 2025). Hence, we initialize most of the workloads with $\approx$ 100 MB to 200 MB of committed memory with the exception of OpenSSL. OpenSSL, within an enclave, allocates memory, performs encryption of a file, and frees the memory. It repeats the process for every new request. Hence, it is initialized with $\approx 460$ MB of committed memory.
Table 2. Description of the workloads.
Workloads Description Input/Setting
BFS (Shun and Blelloch, 2013) Traverse graphs generated by web crawlers. Use breadth-first search. Nodes 150 K Edges 1.9 M
B-Tree (Achermann, 2020a) Create a B-Tree and perform lookup operations on it. Elements: 16 M
HashJoin (Achermann, 2020b) Probe a hash-table (used to implement equijoin in DBs) Lookups: 20 M
Key-Value (Sreekanti et al., 2020) Read and write operations on a key-value store. Elements: 90 M
OpenSSL (OpenSSL, 2019) Encryption-decryption library. File size: 1.1 GB
PageRank (Shun and Blelloch, 2013) Assign ranks to pages based on popularity (used by search engines). Nodes 5000 Edges 12.5 M
SVM (Hearst, 1998) Popular ML algorithm (application: text and hypertext categorization) Rows 10000 Features 128
(a) Total initialization time
(b) Total downtime
(c) Total execution time – single migration
Figure 10. Figure showing end-to-end performance overhead due to migration with MigSGX, TDX, ConstMig-V1 (CM-v1), and ConstMig-V2 (CM-v2). Baseline is SGX + no migration.
6.1. Experimental Setup
We use two 2-socket systems, each with two Intel Xeon Gold 6354 CPUs (72 cores), 1TB of memory, 512GB of SSD, and 1 TB of HDD storage. The system runs the Linux kernel version 5.15. A 1 Gbps network connects the systems. Both the systems support scalable SGX with an EPC of up to 64 GB. We use Intel SGX SDX v2.13 and the SGX driver (Intel, 2019b) for our SGX setup. We use CRIU (Checkpoint and Restore In Userspace) (CRIU Project, 2025) as our migration manager and use LLVM for the compiler pass. Table 2 shows the workloads used to evaluate ConstMig. For TDXevaluation, we use a machine with an Intel Xeon Gold 6526Y CPU, 256GB of main memory, and 2 TB of HDD storage. Our changes are mainly limited to the SGX driver and SDK.
6.2. Evaluation – Single Migration
Here, we evaluate the impact of different migration schemes when an application is migrated just once.
6.2.1. Total Downtime
Figure 10(a) and Figure 10(b) show the total enclave and CVM initialization time and the total downtime experienced by the workloads for a single migration for MigSGX, TDX, ConstMig-V1, and ConstMig-V2. The asynchronous heap migration in ConstMig and TDX, which removes it from the critical path, translates to a significant reduction in the total downtime of an enclave/CVM. ConstMig-V1 reduces the total downtime when compared with MigSGX and ConstMig-V1 by up to 99.4% and 96.2%, respectively. The additional performance gains in ConstMig-V2 are due to a reduction in the time of initialization of a fresh enclave on the destination machine by reducing the committed memory in ConstMig-V2. The impact on the final execution time for the workloads is shown in Figure 10(c). ConstMig-V2 outperforms MigSGX and ConstMig-V1 by an average of $\approx 10\%$ and $\approx 2\%$, respectively. However, the performance difference widens for a long running application with multiple migration (see §6.3).
6.2.2. End-to-End Performance (Single Migration)
Efficient migration also helps in improving the end-to-end performance, as the application is resumed almost instantly on the destination. In terms of the total end-to-end performance of the application being migrated. ConstMig-V1 outperforms MigSGX by up to $\approx$12.61%. ConstMig-V2 outperforms MigSGX by up to $\approx$12.95%. The end-to-end performance gain is due to a quick resumption of the application on the destination machine for ConstMig-V1 and ConstMig-V2. Note that in MigSGX, once data restoration is complete, the application runs without any overhead. However, in ConstMig, even though the application starts responding almost immediately, there are performance overheads due to a background data restore and network page faults. In spite of this, ConstMig outperforms in terms of end-to-end performance.
6.3. At-scale – Multiple Migrations
Here, we measure the performance impact of different migration mechanisms when a long-running B-Tree workload with a working set of 40 GB is migrated multiple times.
Figure 11. Performance overhead in B-Tree with different number of migrations.
As shown in Figure 11, the total execution performance improvement for ConstMig-V2 scales with the number of migrations. For a single migration, ConstMig-V2 outperforms MigSGX, TDX, ConstMig-V1 by $\approx 11\%$, $2.6\%$, and 2.32.3%, respectively. However, for a total of 16 migrations, ConstMig-V2 outperforms MigSGX, TDX, ConstMig-V1 by $\approx 79.6\%$, $23.1\%$, and 48.648.6%, respectively. The scaling in performance improvement is due to the low downtime of the ConstMig-V2 whereas others have a large downtime which adds up with each migration. Note that the migrations are done one after another, there are no “nested” migrations.
6.4. No Migration Performance
Figure 12. Figure showing performance overhead when there is no active migration. FT: Fault Tracker, AT: Access Tracker
Migrations in data centers, if done, are not a frequent operation. Hence, we need to ensure low overhead when the application is not migrating (or is done migrating). Figure 12 shows the overhead experienced by workloads with access tracker and fault tracker. In the access tracker, the additional checks return immediately. There is undoubtedly a small overhead involved; however, a lot of this gets hidden because of the predictability of the branch and the unused resources of out-of-order Intel CPUs. The performance overhead with access tracker is limited to 0.1% to 0.6%. However, with fault tracker, there is no performance overhead as there are no faults generated once the migration is done (or not yet started).
6.5. Deep Dive – Impact on Throughput
We use a reference key-value store (or KVS) workload (Nakashima and Kourai, 2021) to report the observed throughput before, during, and post-migration (MigSGX has done the same). We compare the throughput of ConstMig with state-of-the-art. The KVS workload performs a fixed number of set and get operations in a single thread with a value of size 10 KB. It reports the throughput that it observes. We report the throughput observed before, while, and after the migration. As shown in Figure 13, the throughput goes to zero when the application is frozen. The total downtime is much lower in ConstMig as compared to MigSGX (84% and 96% for ConstMig-V1 and ConstMig-V2, respectively). Furthermore, we observe that post-migration, the observed throughput for ConstMig-V2 is the same as that of native execution, whereas ConstMig-V1 suffers a nominal overhead of $0.1\%$.
Figure 13. Figure showing the drop in the throughput of Key-Value when migrating the application with three different techniques: MigSGX, ConstMig-V1, and ConstMig-V2.
6.6. Network Page Faults (NF)
[…]
Workload BFS HashJoin SVM Rest
#NF 10,064 1,874 12,308 <1000<1000
BFS, a graph search algorithm can reach any node in the tree with few hops. Similarly, in HashJoin, the hash table can be randomly indexed based on the key and the hash value. In SVM, the organization of the trained model is such that it needs to access different addresses scattered in the address space. Hence, they have a high network fault rate (consequence of low spatial locality).
Worst-case scenario: In the worst-case scenario, during migration, every page access on the destination machine causes a network fault. In this case, the total migration time for ConstMig may exceed that of MigSGX. In such extreme and unlikely cases, a stop-and-copy mechanism is suited as the application would anyway be blocked on data to be restored.
7. Related Work
Table 3 presents an overview of the prior work in this area. Gu et al. (Gu et al., 2017) and Yuhala et al. (Yuhala et al., 2021) use the stop-and-copy method to migrate SGX applications from one machine to another. The former focused on the efficient transfer of keys using a remote attestation process, while the latter focused on resuming machine learning applications in case of an application crash. Nakashima and Kourai (Nakashima and Kourai, 2021) and Nakatsuka et al. (Nakatsuka et al., 2022), in their work MigSGX and CTR, respectively, point out that the enclave size can be in GBs, and using the stop-and-copy method might require a large amount of memory. However, their main contribution is in reducing the size of the unsecure buffer required to transfer the enclave data from one machine to another by employing pipelined memory transfer. Their scope is far more limited as compared to our work that looks at full end-to-end migration.
Table 3. A summary of related work
Name SDK Method Memory size
CloudMig  (Gu et al., 2017) Custom Stop-and-Copy $\approx$32MB
Plinius (Yuhala et al., 2021) Simulation Stop-and-Copy & NVMM MBs
MigSGX (Nakashima and Kourai, 2021) Intel Stop-and-Copy & Pipeline GBs
CTR (Nakatsuka et al., 2022) Open Enclave Stop-and-Copy & Pipeline GBs
ConstMig Intel Post-copy GBs
8. Conclusion
In this paper, we present ConstMig– an efficient migration mechanism for large SGX enclaves. We first highlight the scalability limitations of existing approaches, which still rely on the decade-old stop-and-copy migration technique. To address these challenges, we design a secure and efficient save-and-restore mechanism that enables asynchronous migration of enclave data, ensuring constant downtime during migration, regardless of enclave size.
Furthermore, we introduce two novel techniques to prevent the use-before-restore issue in restored enclaves: (i) overloading the SGX page-fault handler and (ii) augmenting each memory access with an additional integrity check. We implement ConstMig and evaluate it using real-world applications. Our results show that ConstMig reduces total downtime by up to 96% for large enclaves compared to the state-of-the-art.
Appendix A Artifact for Dirty Bit Evaluation
The source code for evaluation can be downloaded from here: https://zenodo.org/records/14677356.
A.1. Prerequisite
The system should satisfy the following hardware and software requirements:
  1. (1)
    An Intel SGX machine with Intel SGX SDK and Intel SGX platform software installed.
  2. (2)
    Python
  3. (3)
    Root privileges to access the pagemap.
A.2. Code Structure
The evaluation codebase consists of the following files:
  1. (1)
    py_tracker.cpp: Tracks the dirty bits of a given pid and address ranges using the pagemap interface.
  2. (2)
    normal_app.cpp: Traditional non-SGX application.
  3. (3)
    SecureApp/: A directory that contains an SGX-based application.
To run the py_tracker.cpp file in the background, the user must have root permissions. When executing the normal_app.cpp file, the user should observe that the tracker successfully tracks the dirty memory pages. However, when the secure application is executed, the tracker will be unable to detect dirty memory pages.
A.3. How to run
Please refer to the README file in the code.
A.4. Monitoring PageMap
The code in Listing 2 shows the code that opens the pagemap file from the proc interface for a given pid and reads the dirty bits from the specified address range.
pagemap_path = f’/proc/{pid}/pagemap’
with open(pagemap_path, ’rb’) as f:
f.seek(start_page * 8) # Each entry is 8 bytes
for page in range(start_page, end_page):
entry = f.read(8)
if not entry:
break
pte = struct.unpack(’Q’, entry)[0]
soft_dirty = (pte >> 55) & 1 # Check bit 55
Listing 2: Reading of dirty bit given a pid and address range.